OpenStudy (anonymous):

ALGEBRA!! Factor completely over the complex numbers. x^4+18x^2+81 Any idea how the final answer looked like? (x^2+9)(x^2+9)

OpenStudy (anonymous):

Guys help! @ganeshie8 @Nnesha @Michele_Laino

OpenStudy (anonymous):

@Luigi0210 @sammixboo @welshfella :D

OpenStudy (welshfella):

yea that's correct

OpenStudy (welshfella):

oh you need it over complex numbers

OpenStudy (anonymous):

Yep maybe (x^2+9i)^2 (x^2+9i)^2 ?

OpenStudy (welshfella):

fill in the blanks x^2 + 9 = (x 3i)(x 3i) ?

OpenStudy (anonymous):

(x^2+3i)(x^2+3i) ?

OpenStudy (welshfella):

what signs go in the parentheses?

OpenStudy (anonymous):

Both positives

OpenStudy (welshfella):

no and its x not x^2

OpenStudy (welshfella):

x^2 + 9 = (x + 3i)(x - 3i)

OpenStudy (anonymous):

Oh i get it

OpenStudy (anonymous):

x^4+18x^2+81 this is the that will be factored over the complex numbers

OpenStudy (welshfella):

so complete factors are (x + 3i)(x - 3i)(x + 3i)(x - 3i))

OpenStudy (anonymous):

Is that the same with (x+3i)^2(x-3i)^2 ??

OpenStudy (welshfella):

yes you can write it that way

OpenStudy (welshfella):

are you happy about multiplying these imaginary numbers? -3i * + 3i = -9 * i^2 = -9 * - 1 = +9 because i^2 = -1

OpenStudy (anonymous):



OpenStudy (anonymous):

Looks like this?

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

- where did you get that from? its not x^2 it is x

OpenStudy (welshfella):

If you expanded that the first term would be x^8 not x^4

OpenStudy (welshfella):

thats it

OpenStudy (anonymous):

Thanks! :)

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

Reasons why i love openstudy, and @welshfella Hahahaha

OpenStudy (anonymous):

If i can give you 1000 medals I will